Solve by substitution
6m + n = 11
m - 5n = 38

Answer by Adrish(44)

Solutions:
6m + n = 11...................(1)
m-5n=38.......................(2)
From equation (2) we get,
m= 38+5n
Now substitute the value of m in equation(1) we get,
6(38 + 5n) + n = 11
228 + 30n +n =11
31n + 228 -228= 11-228
31n= -217
n= -217/31
n= -7
substitute the value of n in
m = 38 +5n
m= 38 + 5(-7)
m= 38 - 35
m= 3